OpenStudy (anonymous):

Suppose p varies directly with d, and p = 3 when d = 5. What is the value of d when p = 12?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

'p varies directly with d' so, p=kd from p=3, d=5, first find the value of k

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Whats the next step?

hartnn (hartnn):

now you know 'k' and p=12 so again use p=kd and find 'd'

OpenStudy (anonymous):

I don't know how

hartnn (hartnn):

p=kd 3 = 5k so, k = 3/5

OpenStudy (anonymous):

Continue

hartnn (hartnn):

p=kd 12 = (3/5)d find d from here

OpenStudy (anonymous):

would the answer be 36/5?

OpenStudy (anonymous):

nooooo

OpenStudy (anonymous):

14?

hartnn (hartnn):

d= 12*5/3 now simplify :)

OpenStudy (anonymous):

20!!

hartnn (hartnn):

yes!!